Tuesday, August 19, 2014

8 19 2014 ImageJ Analysis

FTR 230

Hi 60s to low 70s sunny

Participants: Sam Adams and Jake Heare

Sam showed me how to do the ImageJ measurements on the trays today as well as complete several of the trays. The procedure is quite simple.

  1. Open original image in ImageJ
  2. Select line tool
  3. Using on tray size standard (calipers or tile) draw a line
  4. Click M to save after each measurement
  5. repeat this 5 times on the size standard
  6. On each oyster start at the edge of the umbo and draw a line to the furthest point on the bill. 
  7. Click M to save each measure
  8. Proceed measure all oysters in a logical manner from left to right in each quadrant. 
  9. Once all have been measure click Save As in Results window
  10. Save file with the tray label followed by date followed by the word Raw.
  11. Open file in Excell
  12. To get the size conversion you must average of the first 5 "Area" measurements.
  13. Then divide the known size of the Size standard. (10 mm on calipers, 6 inches for tile) by the average you get from the firt 5 measurements.
  14. Using the conversion ratio you can then multiple each ImageJ "Area" measurement by the conversion ratio. 
  15. On the right side of the spreadsheet dedicate 5 columns to the measurements
  16. Column 1 (C1) = Date, C2 = Tray, C3 = Oyster Number, C4 = Area measure, C5 = converted area measurement
  17. Fill in columns with appropriate info
  18. Save file as Excel workbook in eagle/dermochelys/ImageJake with just the tray label and date info
  19. Copy and paste the 5 right side columns into the google sheets master document on Google docs
  20. Repeat process for each image. This should create two files for each image, a raw data and a converted data file. 

Monday, August 18, 2014

8 18 2014 larval counts

Taylor shellfish hatchery, quilcene, wa

Hi 60s to low 80s

Participants brent vadopalas and Jake heare

Went to taylor shellfish to use their z2 culture counter.  took brent and i a couple hours to trouble shoot a tentative protocol for using it. Will have to come back and continue counts for several more days as procedure takes longer than expected. Will provide protocol when it's formally generated in near future. Counts completed today were quite low as compared to numbers provided by dr. Friedman. Brent feels the numbers are accurate.

Friday, August 15, 2014

8 15 14 fidalgo repro check

Anacortes wa
High 50s to low 70s mostly cloudy light rain early dry later
Performed the anesthesia sop with insulation on treatment tub. weather allowed to go without ice blocks or tarp. Found no brooders. Thought that was weird. Checked previous entries and realized I treated the same stack twice in a row due to an error I made last week. Not sure if it really signifies the end of spawning or if I'm really efficient at collecting larvae.

Numbers as follow

Temps in c
Initial.   12
45.       12
1.5.       13
Initial.    12
45.        12
1.5.        12
2.25.      13
Initial.     12
45.         13
1.5.         14

Salinity in ppt
Pretreatment.   25
Treatment.       63
Recovery.        24

Brood collection

Brood.    0
Gaping.   78
Dead.      0
Closed.    25

Brood.    0
Gaping.   83
Dead.      0
Closed.    3

Brood.   0
Gaping.  88
Dead.     0
Closed.   6
Pics have closed on same tile as tube label and calipers. Gapers on all other tiles.

Thursday, August 14, 2014

8 14 14 Oyster Bay Repro Check

Shelton WA

Mid 60's to low 70s mostly cloudy

Performed the anesthesia SOP with insulation on the treatment tub. Weather permitted me to not have to use a tarp or ice blocks. There were no brooders in any population. This is due to the incredibly high innundation of mussels. I cleaned off 95% of the mussel and barnacle mass on the trays and oysters. Hopefully this will allow the animals to feed and reproduce freely. The fouling mass that I removed weighed almost 20 lbs. It was quite impressive. I suggest checking this tray again in September as well as having a general clean up for the other two trays since there would be a lot of mussels on them as well. Surprisingly there were no dead animals as well. I assume the mussels just inhibited reproduction but not feeding. Also relaunched the hobo logger that I thought was possibly damaged. It works fine as of Tuesday (8/12)  so we'll check it in the future.

Numbers as follow.

Temps in C

Initial       13
45           14
1.5          15

Initial       11
45           10
1.5          11
2.25        11

Initial       15
45           17
1.5          18

Salinity in ppt
Pretreatment     27
Treatment         72
Recovery          27

Brood Collection

Brood    0
Gaping   59
Dead      0
Closed   6

Brood    0
Gaping   65
Dead      0
Closed   5

Brood     0
Gaping    52
Dead      0
Closed    0

Pics have tube label and calipers on tile with closed animals. All other tiles are gaping.

This is what a summer of field work does to new boots.

Wednesday, August 13, 2014

8 13 2014 Manchester Repro check

Manchester WA

Hi 60's to Mid 70s Mostly Cloudy

Performed the Anesthesia SOP with insulation and ice blocks on the treatment tub. Weather conditions allowed for the pretreatment and recovery tubs to go uncovered. All populations were brooding this week at Manchester and in almost equal numbers. Had good treatment success due to low boat traffic. Surprised at how active this site has become in supposedly the late part of the season. Definitely want to check this site again around the September full moon.

Numbers as Follow

Temps in C

Initial    9
45        10
1.5       12

Initial     10
45         11
1.5        12
2.25      11

Initial     11
45         12
1.5        12

Salinity in ppt
Pretreatment   29
Treatment       67
Recovery        29

Brood Collection
Brood    3
Gaping   63
Dead      8
Closed   2

#    size    sick
1    31      w
2    24      W
3    22      W

Brood     2
Gaping    59
Dead       8
Closed    0

#    size    sick
1    21      w
2    15      W

Brood    2
Gaping   55
Dead      3
Closed    9

#    size    sick
1    25      W
2    19      W

Pics have brooders between tube label and calipers and closed on same tile. Gapers on all other tiles.

Friday, August 8, 2014

8 8 2014 fidalgo repro check

Anacortes wa

Mid 60s to high 70s sunny

Performed the anesthesia sop with insulation on treatment tub and a tarp on the Pretreatment and recovery tubs. Treatment was effective as normal so it seemed to be a fluke the last couple of weeks. Also had pretty equal brooding across all populations. Finally the northern population had the most number of brooders. I think we should check these animals again around the full moon in September to see if they are still spawning.

Numbers as follow :

Temps in c
Initial.   11
45.      13
1.5.       15

Initial.     13
45.        14
1.5.         13
2.25.       14

Initial.     11
45.        12
1.5.         15

Salinity in ppt
Pretreatment.       27
Treatment.           56
Recovery.            28

Brood collection
Brood.      2
Gaping.     70
Dead.        0
Closed.      22

#   size.  Sick
1.     32.    Grey
2.     30.    W

Brood.       3
Gaping.      67
Dead.         0
Closed.      15

#     size.     Sick
1.      30.       W
2.     29.        W
3.      20.      W

Brood.     5
Gaping.    89
Dead.       0
Closed.    10

#   size.   Sick
1.     19.      W
2.    34.      W
3.    35.      W
4.    32.      W
5.    29.      W

Pics have brooders between Tube and calipers. Closed on same tile. Gaping on all other tiles.

Thursday, August 7, 2014

8 7 2014 Oyster Bay repro check

Shelton wa

Mid 60s to low 80s sunny

Performed the anesthesia sop with insulation on treatment tub and ice blocks on treatment and tarp cover on the Pretreatment and recovery tubs. It appears that the mussels have a significant effect on brooding as today's trays were less cover or not covered at all compared to last weeks. I saw a substantial number of brooders in the southern pop and a few in the northern and dabob pop.  We could almost assume that the other trays should be brooding in the same number as the "clean trays". We'll find out next week if the cleaning job I did last week helped out.

Numbers as follow:

Temps in c
Initial.    15
45.       15
1.5.       18

Initial.     13
45.        14
1.5.         13
2.25.      12

Initial.      14
45.         15
1.5.          18

Salinity in ppt
Pretreatment.        25
Treatment.            55
Recovery.             27

Brood collection

Brood.    2
Gaping.   80
Dead.      0
Closed.    6

#    size.    Sick
1.     29.       W
2.    31.        W

Brood.      10
Gaping.     80
Dead.        0
Closed.     1

#     size.    Sick
1.      30.      W
2.     34.       W
3.     31.        Grey
4.     33.       W
5.    35.       W
6.    27.        W
7.    29.         W
8.    26.        Grey
9.    25.         W
10.   25.        W

Brood.    3
Gaping.   71
Dead.     0
Closed.   7

#   size.   Sick
1.    28.      Grey
2.   30.      W
3.   37.       W

Pics have brooders between tube and calipers. Closed on same tile gapers on other three tiles.

Wednesday, August 6, 2014

8 6 2014 Manchester Repro check

Manchester WA

Hi 60s to Low 80s

Performed the anesthesia SOP with insulation and ice blocks on the treatment tub and tarp on the pretreatment and recovery tubs. Forgot my work box when I was packing yesterday and had to borrow a digital thermometer, ruler, refractometer, and notebook from Doug and Andy at the NOAA facility. The thermometer read in Fahrenheit and the refractometer hadn't been calibrated recently. The readings from both seemed accurate enough though the Temp readings will need to be converted to Celsius at some point. As for spawning, the animals at Manchester seem to be increasing their spawning efforts as the next tidal cycle approaches. I had 4 brooders in the southern pop and 2 in the dabob pop, all of pretty good quality. It wouldn't surprise me with the temperatures that if the current temperatures remain through the middle of September these animals might continue spawning until they are forced to stop by temp influence.

Numbers as Follow:

Temps in F
Initial    64
45        66
1.5       67

Initial    68
45        69
1.5       70
2.25     69

Initial   61
45       63
1.5      64

Salinity in ppt
Pretreatment      33
Treatment          65
Recovery           34

Brood Collection

Brood   2
Gaping  77
Dead     12
Closed  5

#    Size    Sick
1     24      W
2     24      W

Brood     0
Gaping    45
Dead       4
Closed    5

Brood     4
Gaping    50
Dead       11
Closed    7

#    size    sick
1    24      W
2    22      W
3    25      W
4    25      Grey

Pics are of all animals. Brooders are between purple ruler and tube label. Closed are on the same tile as brooders. Gaping animals are on all other tiles. Some crossover due to adhered animals on tiles.

Friday, August 1, 2014

8 1 2014 Fidalgo Repro Check

Anacortes WA

Mid 60s to high 70s. Sunny until 2 pm then partly cloudy

Performed the anesthesia SOP with insulation on the treatment tub and ice blocks in the Treatment and recovery tubs to stabilize temperature fluctuations. Had a really poor treatment of the north sound tray and a less than great treatment of the south pop tray. Less than a third of the animals were gaping in north pop and only a little over half in the South pop were open. The Dabob tray was over 95% treated. I'm not sure we missed any brooders though as South sound only had a few and Dabob had none. I'm not sure why the animals didn't open but I assume that it might be due to the fact that they were quickly sprayed with fresh water to help knock off some of the more intense sediment deposits on them. This great to know for next time so I don't accidentally deter treatment. It seems like spawning season is winding down in Fidalgo. I'll be interested to see what the next tidal cycle does. Also yesterday I had a problem with the Hobo Shuttle. It would not read the logger pendant. I brought the logger pendant and shuttle with me to make sure neither of them was screwed up. Today I retried the pendant with the shuttle and the data offload properly. I think what happened yesterday was the shuttle go hot sitting in the sun in my tool box. The air was pretty stagnant and it got much hotter than expected. I'll double check everything on Monday to make sure the file and shuttle are in working order.

Numbers as Follow:

Temps in C

Pretreatment * Lacking the tarp and with little breeze, the trays changed temps faster than expected.
Initial  10
45      15
1.5     15

Initial     13
45         13
1.5        14
2.25      14

Initial    15
45        15
1.5       15

Salinity in ppt
Pretreatment    27
Treatment        60
Recovery         28

Brood Collection

Brood    0
Gaping   28
Dead      0
Closed   68

Brood     3
Gaping    45
Dead       0
Closed    38

#    size    sick
1    30      W
2    35      W
3    33      W

Brood   0
Gaping  84
Dead     0
Closed  2

As for Pics. There were so few treated in the North pop that the tube label and calipers are on the tile with the gapers and the rest of the tiles are closed. The South pop tray is like normal with the brooders between the tube label and calipers. The Dabob tray had so many gapers that they are on all tiles but the closed guys are above the tube and calipers.

My face when treatments don't work.